Exercise 5S-9 Equivalent Units; Equivalent Units of Production; Assigning Costs-FIFO Method [LO5S-6, LO5S-7, LO5S-8]

Jarvene Corporation uses the FIFO method in its process costing system. The following data are for the most recent month of operations in one of the company’s processing departments:

Units in beginning inventory 410
Units started into production 4,280
Units in ending inventory 290
Units transferred to the next department 4,400
Materials Conversion
Percentage completion of beginning inventory 70 % 30 %
Percentage completion of ending inventory 80 % 40 %

The cost of beginning inventory according to the company’s costing system was $7,849 of which $4,841 was for materials and the remainder was for conversion cost. The costs added during the month amounted to $179,249. The costs per equivalent unit for the month were:

Materials Conversion
Cost per equivalent unit $18.00 $23.00

Required:

1. Compute the total cost per equivalent unit for the month.

2. Compute the equivalent units of material and conversion in the ending inventory.

3. Compute the equivalent units of material and conversion that were required to complete the beginning inventory.

4. Compute the number of units started and completed during the month.

5. Compute the cost of ending work in process inventory for materials, conversion, and in total for the month.

6. Compute the cost of the units transferred to the next department for materials, conversion, and in total for the month.
Answer #1

1) Total cost per equivalent unit = 18+23 = 41

2) Equivalent unit of material and conversion

Material Conversion
Ending inventory 290*80% = 232 290*40% = 116

3) Equivalent unit of material and conversion

Material Conversion
Beginning inventory 410*30% = 123 410*70% = 287

4) Unit started and completed = 4400-410 = 3990

5) Cost of ending WIP

Material Conversion Total
Cost of ending WIP 232*18 = 4176 116*23 = 2668 6844

6) Cost of units transferred to next department

Material Conversion Total
Cost of Unit transferred out (123+3990)*18 = 74034+4841 = 78875 (287+3990)*23 = 98371+3008 = 101379 180254